Solve
x^2y''- 2xy'+2(1+x^2)y=0 , x>0
1
Expert's answer
2020-12-14T10:21:52-0500

"x^2y''-2xy'+2(1+x^2)y=0, x>0\\"

Divide through by "x^2"

"y''-\frac{2}{x}y'+\frac{2(1+x^2)}{x^2}y=0\\"

For the equation of the form "y''+f(x)y'+g(x)y=F(x)." The substitution "y=v.e^{- \int \frac{f(x)}{2}dx}" is used.


So,

"y=v.e^{- \int \frac{-2}{2x}dx}\\
y=v.e^{\int\frac{1}{x}dx}\\
y=v.e^{\ln x}\\
y=vx\\
y'=v+xv'\\
y''=2v'+xv''"

Rewrite the equation

"(2v'+xv'')-\frac{2}{x}(v+xv')+\frac{2(1+x^2)}{x^2}vx=0\\
xv''+2v'-\frac{2v}{x}-2v'+\frac{2(1+x^2)}{x}v=0\\
xv''-\frac{2}{x}v+\frac{2(1+x^2)}{x}v=0\\
xv''+\frac{2}{x}v(-1+1+x^2)=0\\
xv''+\frac{2}{x}v(x^2)=0\\
xv''+2xv=0"

divide through by "x"

"v''+2v=0" .


The associated auxilliary equation to this is ;

"m^2+2m=0\\
m(m+2)=0\\
m=0,m=-2"

So,

"v=A_1e^{0.x}+A_2e^{-2.x}\\
v=A_1+A_2e^{-2x}"


But "v=\frac{y}{x}"

"\frac{y}{x}=A_1+A_2e^{-2x}\\
y=x(A_1+A_2e^{-2x}) ,x>0"